3. Find the angle of the largest right circular cone which can be inscribed in a sphere of





radius 9 inches.





4. A statue 10 feet high is standing on a base 13 feet high. If an observer’s eye is 5 feet





above the ground, how far should he stand from the base in order that the angle





between his lines of sight to the top and bottom of the statue is a maximum. (How far





should he stand to get the best view of the statue.





5. A steel girder 27 feet long is to be moved on rollers along a passageway 8 feet in





width and into a corridor at right angles to the passageway. If the horizontal width of





the girder is neglected, how wide must the corridor be in order that the girder can go





around the corner?

1
Expert's answer
2021-11-29T00:17:34-0500

4.


If the angles to the bottom and top are a and b, then from a distance x we have


"tan(a) = \frac{8}{x}"


"tan(b) = \frac{18}{x}"


We want the angle "c=b-a", so


"c = arctan(\frac{18}{x})-arctan(\frac{8}{x})"


"\frac{dc}{dx} = \frac{10(144-x^2)}{((x^2+8^2)(x^2+18^2))}"


"\frac{dc}{dx}=0" at


"x=12"